How do you solve for:
x+12y=6
x-1/3y=4

Answer by stanbon(75887)

solve for:
x+12y=6
x-1/3y=4
---------------
Subtract the 2nd from the 1st equation to get:
12 1/3 y = 2
---
(37/3)y = 2
y = 2(3/37)
y = 6/37
=============
Substitue for y in (x +12y = 6) and solve for "x":
x + 12(6/37) = 6
---
x + 72/37 = 6
x = 150/37
=================
Cheers,
Stan H.
=============